At a certain temperature, 0.820 mol SO3 is placed in a 4.00 L container. 2SO3(g)↽−−⇀2SO2(g)+O2(g) At equilibrium, 0.130 mol O2 is present. Calculate ?c.
Answer #1

Answer:

Explanation:

Kc is defined as the amounts of products and reactants present at equilibrium in a reversible chemical reaction at a given temperature

Step 1: Write the balanced the chemical equation

2SO3 (g) <--------------> 2SO2 (g)+O2(g)

Step 2: Calculate the Concentrations by dividing the given moles by volume of container.

Initial SO3 = (0.820 moles / 4 litre ) = 0.205 M

as well the concentration of O2 at equilibrium = ( 0.130 moles / 4 L ) = 0.0325 M

Step 3: Write the ICE table

  2SO3 (g) <--------------> 2SO2 (g) + O2(g)

Initial amount 0.205 0 0

Change amount -2x +2x +x

Equilibrium amount 0.205-2x 2x x

Since the value of O2 at equlibrium is given [O2] = 0.0325 M = [x]

hence value of x = 0.0325 M

Hence all value at equilibrium is

[SO3] = 0.205 -2x = 0.205 -(2 × 0.0325 ) =0.14

[SO2] = 2x = 2 × 0.0325 = 0.065

[O2] = x = 0.0325

Step 4: Calculate the Kc

Kc = [SO2]2 × [O2] / [SO3]2 = ( 0.065)2×(0.0325) / (0.14)2 = 7 ×10-3